Last year Rachel invested money in two accounts. The first account had an interest rate of 6% and the second account had an interest rate of 4%. If she invested $300 more in the first account than the second and her total interest income was $608, how much did she invest at each rate?
Step 1 of 3 : Complete the following table by entering the missing values. Enter the rates in decimal for

Answer by stanbon(75887)

Last year Rachel invested money in two accounts. The first account had an interest rate of 6% and the second account had an interest rate of 4%. If she invested $300 more in the first account than the second and her total interest income was $608, how much did she invest at each rate?
----
Equations:
s = f + 300 dollars
0.06s + 0.04f = 608 dollars
-------------------------------------
Modify for elimination::
6s - 6f = 6*300
6s + 4f = 100*608
-----------------------
Subtract and solve for "f"::
10f = 59000
-----
f = $5900.00 (amt invested at 4%)
s = f+300 = $59,300 (amt invested at 6%)
--------------
Cheers,
Stan H.
--------------